# Performance study of 4-MU-loaded water for Cherenkov light detection

### v1

@@ 1 Introduction
Water Cherenkov detectors are widely used in particle and astroparticle physics for their scalability, directional sensitivity, and relatively low cost. Their applications range from neutrino detection to cosmic-ray veto systems  [FUKUDA2003418, Knapp:2009zz, Geis:2018com]. In such detectors, charged particles traveling faster than the speed of light in water emit Cherenkov photons, which are subsequently detected by photomultiplier tubes (PMTs) lining the detector volume. However, the number of detectable Cherenkov photons is often limited by the wavelength-dependent sensitivity of PMTs and the absorption properties of water. Improving the light yield and collection efficiency can directly enhance the detector performance, particularly for low-energy or rare-event signals.
The AMoRE-II (Advanced Mo-based Rare process Experiment) project aims to search for neutrinoless double beta decay using cryogenic detectors operated in a deep underground laboratory. To suppress muon-induced backgrounds, the experiment employs a large water Cherenkov detector as an active muon veto system, filled with about 60 tons of deionized water and instrumented with 48 PMTs to detect Cherenkov light produced by passing muons [AMoRE:2024tjb]. Enhancing the photon detection efficiency of this veto detector is essential for improving the background rejection power and ensuring the sensitivity goals of the experiment.
One approach to increase the number of detectable photons in water Cherenkov detectors is to use wavelength shifters (WLS) that absorb ultraviolet Cherenkov light and re-emit it at longer wavelengths, where PMTs are more efficient [Dai:2008cp]. Among various candidates, 4-methylumbelliferone (4-MU) is known for its strong absorption in the UV region and efficient fluorescence in the blue range, making it a suitable WLS additive in water-based systems [Sweany:2011qh].
@@ 2 Optical Properties of 4-MU
To evaluate the optical behavior of 4-MU [AlfaAesarA10337, ChemicalBook4MU, Cayman4MU] in water, we measured its absorbance and fluorescence emission spectra. These measurements are relevant for assessing the potential of 4-MU as a wavelength-shifting material in water Cherenkov detectors.

### v2

@@ Abstract
(Accepted for publication in the Journal of Instrumentation (JINST), 27 January 2026.)
@@ 1 Introduction
Water Cherenkov detectors are widely used in particle and astroparticle physics for their scalability, directional sensitivity, and relatively low cost. Their applications range from neutrino detection to cosmic-ray veto systems  [5, 7, 6]. In such detectors, charged particles traveling faster than the speed of light in water emit Cherenkov photons, which are subsequently detected by photomultiplier tubes (PMTs) lining the detector volume. However, the number of detectable Cherenkov photons is often limited by the wavelength-dependent sensitivity of PMTs and the absorption properties of water. Improving the light yield and collection efficiency can directly enhance the detector performance, particularly for low-energy or rare-event signals.
The AMoRE-II (Advanced Mo-based Rare process Experiment) project aims to search for neutrinoless double beta decay using cryogenic detectors operated in a deep underground laboratory. To suppress muon-induced backgrounds, the experiment employs a large water Cherenkov detector as an active muon veto system, filled with about 60 tons of deionized water and instrumented with 48 PMTs to detect Cherenkov light produced by passing muons [1]. Enhancing the photon detection efficiency of this veto detector is essential for improving the background rejection power and ensuring the sensitivity goals of the experiment.
One approach to increase the number of detectable photons in water Cherenkov detectors is to use wavelength shifters (WLS) that absorb ultraviolet Cherenkov light and re-emit it at longer wavelengths, where PMTs are more efficient [4]. Among various candidates, 4-methylumbelliferone (4-MU) is known for its strong absorption in the UV region and efficient fluorescence in the blue range, making it a suitable WLS additive in water-based systems [8].
@@ 2 Optical Properties of 4-MU
To evaluate the optical behavior of 4-MU [9, 3, 2] in water, we measured its absorbance and fluorescence emission spectra. These measurements are relevant for assessing the potential of 4-MU as a wavelength-shifting material in water Cherenkov detectors.